You are playing a game. You are picking one ball out of a bowl with 15 balls in it (assume each ball has the same probability to be picked) – the following balls are in the bowl:




5 red balls (you win/lose 0 Euro)




9 green balls (you lose 1 Euro)




1 blue ball (you win 10 Euro)




How much money do you win or lose in the long run per trial? (1p.)

1
Expert's answer
2021-11-18T07:09:47-0500

you win or lose in the long run per trial:


"E(X)=\sum p_ix_i=-\frac{1}{C^9_{15}}+\frac{1}{15}\cdot 10=-\frac{6!9!}{15!}+2/3=0.6667-0.0002=0.6665" Euro